OpenStudy (anonymous):

what is sin 0 in the triangle MEDAL AND FAN

OpenStudy (anonymous):

5/13

OpenStudy (anonymous):

how do i do that ??

OpenStudy (amorfide):

sohcahtoa sin so you use oh o is the side length opposite the angle h is always side length opposite the right angle sinx=o/h substitute your values sinx=10/26

OpenStudy (anonymous):

lol thanks @amorfide 5/13

OpenStudy (amorfide):

no problem